OpenStudy (anonymous):

how to clear then substitution x/2- y/3=5/6 x/5-y/4=29/10

OpenStudy (anonymous):

multiply the first by 6 the second by 20

OpenStudy (asnaseer):

in questions like this, it is usually best to first get rid of the fractions

OpenStudy (asnaseer):

you can do that in the first equation by multiplying it by \(2*3=6\) - that will remove both the denominators 2 and 3

OpenStudy (asnaseer):

similarly, multiply the second equation by \(5*4=20\) - which will get rid of both the denominators 5 and 4

OpenStudy (asnaseer):

write down what you get after these steps

OpenStudy (anonymous):

6x-6y=30 5x-4y=20

OpenStudy (asnaseer):

that does not look correct

OpenStudy (asnaseer):

let me do the first one for you - and then you can reattempt the second one

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):